# Self-Reproduction and Evolution in Cellular Automata: 25 Years after Evoloops

### v1

@@ 4 Future Challenges and Open Problems
These recent developments that utilize continuous cellular automata are very promising and expected to produce further advances in self-replication and evolution research. Meanwhile, many key questions still need to be addressed, and here we discuss only a few.
First, there are many concrete challenges and open problems in self-reproduction and evolution for von Neumann’s self-reproducing universal constructors:
@@
Second, nearly all the above evolutionary systems built within spatially distributed media exhibited the eventual dominance by one or a few most successful species in the long run, and it is still unclear what kind of generalizable principles or mechanisms are available to prevent the evolving ecosystem of self-replicators from falling into such pseudo-equilibrium states. It has been suggested that dynamic environments [45, 53, 2, 1] are the key to addressing this issue, although they may not work for indefinitely long terms [54].
Third, the recent approach to assign model parameters to local regions/agents deliberately avoids explicit representation of such genetic information in the space, in contrast to real biological systems and earlier self-replication models [68, 25, 48, 49, 50] where genetic instructions were explicitly represented in space. It is not well understood how these two approaches differ regarding the open-endedness and creativity of their evolution.
Fourth, all the models reviewed so far relied on either logically designed mechanisms written in discrete state-transition rules or dynamically generated quasi-stable patterns in continuous space, but neither would capture the autopoietic nature of real biological systems. There is another major body of literature on computational autopoietic models [30, 37, 22, 61, 57, 58], and some of them even demonstrated self-reproduction of autopoietic structures [58]. However, it remains unclear how one could integrate autopoietic dynamics into existing cellular automata models of evolving self-replicating patterns.

### v2

@@ 4 Future Challenges and Open Problems
These recent developments that utilize continuous cellular automata are very promising and expected to produce further advances in self-reproduction and evolution research. Meanwhile, many key questions still need to be addressed, and here we discuss only a few.
First, there are many concrete challenges and open problems in self-reproduction and evolution for von Neumann’s self-reproducing universal constructors [73, 77]. One group concerns extensions of von Neumann’s original 29-state cellular automaton model of self-reproduction (using strategy Mech. 2):
@@
A further set of hard challenges refers to self-reproducers going beyond the methods of the original von Neumann self-reproducer (that was completed by A.W. Burks and others):
@@
Second, nearly all the above evolutionary systems built within spatially distributed media exhibited the eventual dominance by one or a few most successful species in the long run, and it is still unclear what kind of generalizable principles or mechanisms are available to prevent the evolving ecosystem of self-replicators from falling into such pseudo-equilibrium states. It has been suggested that dynamic environments [51, 59, 2, 1] are the key to addressing this issue, although they may not work for indefinitely long terms [60]. Overcoming this empirical limitation is a necessary and critical step towards implementing open-ended self-reproducing and evolving systems within cellular automata and other similar spatially distributed computational media.
Third, the recent approach to assign model parameters to local regions/agents deliberately avoids explicit representation of such genetic information in the space, in contrast to real biological systems and earlier self-replication models [74, 27, 54, 55, 56] where genetic instructions were explicitly represented in space. It is not well understood how these two approaches differ regarding the open-endedness and creativity of their evolution. They also differ significantly with regard to the computational capability built in the environment and in the “laws of physics”. Gaining an in-depth understanding of the pros and cons of these two approaches will greatly inform us about the design choices for future models of self-reproduction and evolution in cellular automata.
Fourth, all the models reviewed so far relied on either logically designed mechanisms written in discrete state-transition rules or dynamically generated quasi-stable patterns in continuous space, but neither would capture the autopoietic nature of real biological systems. There is another major body of literature on computational autopoietic models [34, 41, 23, 67, 63, 64, 25, 30], and some of them even demonstrated self-replication of autopoietic structures [64, 30]. However, it remains unclear how one could integrate autopoietic dynamics into existing cellular automata models of evolving self-reproducing patterns. This is a potentially promising yet largely under-explored subject of study on cellular automata-based self-reproduction and evolution, which might eventually lead to a completely brand-new approach.